OpenStudy (anonymous):

2^(x+3)=5^x Help solve for x

OpenStudy (anonymous):

can someone show me the steps

OpenStudy (anonymous):

@zepdrix pls, i dont get it!

geerky42 (geerky42):

Try take log of both sides.

geerky42 (geerky42):

Like \(\ln(2^{x+3})=\ln(5^x)\). Then you can use this logarithm property \(\log(a^b) = b\log(a)\)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is a raft of algebra after the first step

OpenStudy (misty1212):

\[2^{x+3}=5^x\] first step is \[(x+3)\ln(2)=x\ln(5)\] then solve for \(x\) actually not too much algebra in this one

OpenStudy (misty1212):

distribute and get \[\ln(2)x+3\ln(2)=x\ln(5)\] put all the \(x\) terms on the right and get \[3\ln(2)=\ln(5)x-\ln(2)x\] then factor out the \(x\) \[3\ln(2)=(\ln(5)-\ln(2))x\] and finally divide